OpenStudy (anonymous):

find all points having an x-coordinate of 1 whose distance from the point (-2 -3) is 5

OpenStudy (anonymous):

\[ (1+2)^2 +(y+3)^2 =25 \\ 9+(y+3)^2 =25\\ (y+3)^2 =16\\ y+3= \pm 4 \] Can you finish it?

OpenStudy (anonymous):

no :-(

OpenStudy (anonymous):

\[ y+ 3 = 4\\ y=1\\ y+3 = -4\\ y=-7 \] The points are (1,1) and ( 1,-7)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i ahve another question if you don't mind...

OpenStudy (anonymous):

ok thanks